A spherical ball has a radius of 0.2 m and electric charge is 20uc in its center.Calculate the electric flux pass through the spherical ball.




1
Expert's answer
2022-04-11T16:55:42-0400

"E=\frac{kq}{r^2}=4.5~\frac{MN}C."